OpenStudy (anonymous):

solve the following equation x^2-4x-3=0

OpenStudy (anonymous):

use sridharacharyas formula

OpenStudy (anonymous):

what's that formula?

OpenStudy (anonymous):

i have no idea lol

OpenStudy (anonymous):

search on google

OpenStudy (anonymous):

come to this site for help and you send me to another lol

OpenStudy (anonymous):

\[(4\pm \sqrt{16-12})\div2\]

hartnn (hartnn):

it means quadratic formula: Compare your quadratic equation with \(ax^2+bx+c=0\) find a,b,c then the two roots of x are: \(\huge{x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}}\)

OpenStudy (anonymous):

@Joseph91 right

OpenStudy (anonymous):

Either X=1or X=3

hartnn (hartnn):

x is not 1

hartnn (hartnn):

neither 3, lol!

OpenStudy (anonymous):

you can use the modes in calculator

hartnn (hartnn):

@rodneyvs2000 use the formula i gave, u will get your roots.

OpenStudy (anonymous):

thanks all

hartnn (hartnn):

but did u get your answer as \(\huge2\pm\sqrt7\)